OpenStudy (zab505):

Suppose that a restaurant chain claims that its bottles of ketchup contain 24 ounces of ketchup on average, with a standard deviation of 0.8 ounces. If you took a sample of 49 bottles of ketchup, what would be the approximate 95% confidence interval for the mean number of ounces of ketchup per bottle in the sample? A. 24 ± 0.029 B. 24 ± 0.057 C. 24 ± 0.114 D. 24 ± 0.229

OpenStudy (anonymous):

a

OpenStudy (anonymous):

Elaborate @cram ?

OpenStudy (anonymous):

okay if you take 0.8 and times it by 49 /95% there is the answer

OpenStudy (anonymous):

The confidence interval for a \((1-\alpha)100\%\) confidence level is given by\[\left(\bar{x}-Z_{\alpha/2}\frac{\sigma}{\sqrt n},~\bar{x}+Z_{\alpha/2}\frac{\sigma}{\sqrt n}\right)\] with \(\bar{x}\) being the sample mean (24), \(Z_{\alpha/2}\) the critical value (1.96 for 95% confidence), \(\sigma\) the standard deviation (0.8), and \(n\) the sample size (49). \[\left(24-1.96\frac{0.8}{\sqrt {49}},~24+1.96\frac{0.8}{\sqrt {49}}\right)\] Basically it's a matter of simplifying the margin of error term (after the \(\pm\)). \[1.96\frac{0.8}{7}\approx0.224\] which makes D the closest answer.